OpenStudy (anonymous):

Passes through 2,6 and is perpendicular to y=3x-6

OpenStudy (anonymous):

@SolomonZelman I have no clue on how to do perpendicular ones o; I only have 2 of them to do though

OpenStudy (michele_laino):

please note that the slope m of your line is : \[m=-\frac{ 1 }{ 3 }\] because product of the slopes of perpendicular lines has to be equal to 1 so,please apply the subsequent equation: \[y-y _{0}=m(x-x _{0})\] where (x_0,y_0) is your point and m=-1/3

OpenStudy (michele_laino):

oop the product of the slopes of perpendicular lines has to be equal to -1

OpenStudy (anonymous):

o;

OpenStudy (dtan5457):

Yes. A perpendicular slope of any slope is flipped and negated. If the original slope was -1/3, your perpendicular slope would be 3/1.

OpenStudy (anonymous):

1/12 is the answer

OpenStudy (dtan5457):

Do you know where to go from there?

OpenStudy (anonymous):

thats what i got

OpenStudy (anonymous):

y=3x+1/12

OpenStudy (michele_laino):

@dtan5457 please note that the original slope is 3

OpenStudy (dtan5457):

@Michele_Laino I am aware. I was just setting an example.

OpenStudy (michele_laino):

oops.. sorry @dtan5457

OpenStudy (anonymous):

thanks to both of u :D

OpenStudy (anonymous):

you guys have time for 1 more?

OpenStudy (michele_laino):

thank you! @ayeitsJenni.

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

passes through -1,5 and is perpendicular to y=-5/2x+2

OpenStudy (anonymous):

im not good with fractions.

OpenStudy (michele_laino):

please, note that if I have this fraction: 5/7 the inverse fraction is: 7/5

OpenStudy (michele_laino):

so what in the inverse fraction of 5/2?

OpenStudy (anonymous):

don't you only inverse at the end? when you find the slope which is b..

OpenStudy (anonymous):

2/5

OpenStudy (michele_laino):

ok! now please change its sign and you will get your slope!

OpenStudy (anonymous):

-2/5?

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

now, please apply the formula above, where, this time, (x_0,y_0) is the point (-1,5)

OpenStudy (michele_laino):

and m=-2/5

OpenStudy (anonymous):

I don't really get none of this >_<

OpenStudy (anonymous):

don't you use y=mx+b?

OpenStudy (michele_laino):

please, you have to apply this formula: \[y-y _{0}=m(x-x _{0)}\] where m=-2/5 as you wote, and (x_0,y_0)=(-1,5)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so its y - 5 = -2/5(x--1)

OpenStudy (michele_laino):

ok! the parentheses, is: \[(x-(-1))=(x+1)\]

OpenStudy (anonymous):

y-5 = -2/5 (x+1)

OpenStudy (anonymous):

thats the equation now?

OpenStudy (michele_laino):

yes! it is!

OpenStudy (anonymous):

oh awesome

OpenStudy (anonymous):

what do i do now

OpenStudy (michele_laino):

you have to perform multiplication, namely: \[-\frac{ 2 }{ 5 }(x+1)=...\]

OpenStudy (anonymous):

-2/5x + -2/5

OpenStudy (anonymous):

is that it?

OpenStudy (michele_laino):

yes! please note that: \[+-\frac{ 2 }{ 5 }=-\frac{ 2 }{ 5 }\]

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

ok! now you have this equation: \[y-5=-\frac{ 2 }{ 5 }x-\frac{ 2 }{ 5 }\] now, please add to both sides 5

OpenStudy (anonymous):

y = -2/5x-2/5+5

OpenStudy (michele_laino):

ok! now you have to perform this addition, please: \[-\frac{ 2 }{ 5 }+5=...\]

OpenStudy (anonymous):

kinda finding it hard to do that

OpenStudy (michele_laino):

note that the last common multiple is 5, so: \[-\frac{ 2 }{ 5 }+5=\frac{ -2+5*5 }{ 5 }=\frac{ -2+25 }{ 5 }=\frac{ 23 }{ 5 }\]

OpenStudy (michele_laino):

is it clear, please?

OpenStudy (anonymous):

how do I write down the answer as a fraction, and yes...

OpenStudy (michele_laino):

ok! now please write your equation

OpenStudy (anonymous):

i mean, an equation

OpenStudy (anonymous):

How do i write the equation o;

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

y=-2/5x+23/5?

OpenStudy (michele_laino):

well done! you are welcome!!!!! :)

OpenStudy (anonymous):

thanks

OpenStudy (michele_laino):

thanks!